Solve by completing the square.
3x^2-9x-31=0

Answer by stanbon(75887)

Solve by completing the square.
3x^2-9x-31=0
3(x^2-3x+(3/2)^2=31+3(3/2)^2
3(x+(3/2))^2 = 151/4
(x+(3/2))^2=151/12
x= (-3/2) + (sqrt(151/12)) or x=(-3/2)-(sqrt(151/12))
x=-2.0473 or x=-5.047
=========================
Cheers,
Stan H.